### v1

# ChEX: Interactive Localization and Region Description in Chest X-rays

Abstract
Report generation models offer fine-grained textual interpretations of medical images like chest X-rays, yet they often lack interactivity (i.e. the ability to steer the generation process through user queries) and localized interpretability (i.e. visually grounding their predictions), which we deem essential for future adoption in clinical practice. While there have been efforts to tackle these issues, they are either limited in their interactivity by not supporting textual queries or fail to also offer localized interpretability. Therefore, we propose a novel multitask architecture and training paradigm integrating textual prompts and bounding boxes for diverse aspects like anatomical regions and pathologies. We call this approach the Chest X-Ray Explainer (ChEX). Evaluations across a heterogeneous set of 9 chest X-ray tasks, including localized image interpretation and report generation, showcase its competitiveness with SOTA models while additional analysis demonstrates ChEX’s interactive capabilities.
Keywords:
Vision-Language Modeling Localized Report Generation1 Introduction
The automatic interpretation of medical images, such as chest X-rays, holds immense promise for enhancing healthcare. Report generation models enable detailed textual interpretations beyond the capabilities of traditional image classification techniques alone. Despite significant progress in enhancing their accuracy[26], incorporating them in clinical practice remains a formidable challenge. A primary concern stems from the lack of transparency and interpretability surrounding the decision-making processes of these models, which poses obstacles for medical professionals seeking to validate their predictions, thereby hindering widespread adoption[48, 13]. Additionally, the non-interactive nature of most current models introduces risks in cases of inaccurate predictions, while interactive involvement in the generation process, e.g. through user prompts, may encourage the user to manually intervene in such cases.
Therefore, we advocate for two important aspects that can enhance the utility of these models in clinical practice: (localized) interpretability and interactivity. One pioneering effort in this direction is the work of Tanida et al. [69]. By incorporating bounding boxes of anatomical regions into the report generation process, their model offers enhanced interpretability. Furthermore, it supports bounding boxes as input queries, facilitating user interactivity. However, the model exhibits two major limitations: it lacks support for textual user prompts and focuses exclusively on anatomical regions during both training and inference. This narrow focus can lead to suboptimal predictions for other aspects, such as pathologies. In contrast, models like RaDialog [56], Med-PaLM M [71], and OmniFM-DR [82] support interactivity through textual prompts but do not provide bounding boxes for their textual answers or support them as queries.
In this work, we address these limitations by proposing a novel multitask architecture and training paradigm. Our approach integrates textual prompts and bounding boxes for various aspects, including anatomical regions, pathologies, and report sentences. We call this approach the Chest X-Ray Explainer (ChEX). As illustrated in Figure 1, ChEX can be queried using textual prompts or user-defined bounding boxes and predicts individual descriptions for each specified region or aspect. For textual queries, the predictions are supplemented by bounding boxes to localize relevant regions for each answer. Therefore, ChEX offers a unique combination of interactivity – reacting to user prompts – and interpretability – visually grounding the answers – not provided by other report generation models. Furthermore, it additionally supports localized tasks beyond report generation (RG), including pathology (object) detection (OD), sentence grounding (SG), region classification (RC), and region explanation (RE).
Our contributions are as follows:
We propose ChEX, an interactive and interpretable model for predicting visually grounded textual descriptions of chest X-rays based on user queries.
We propose a multitask training paradigm, enabling ChEX to be jointly trained with diverse types of supervision from different datasets.
We evaluate ChEX across 9 diverse chest X-ray tasks, spanning localized image interpretation and report generation functionalities. ChEX demonstrates competitive performance with specialized and general state-of-the-art (SOTA) models despite being significantly smaller than some of these models, which have up to 80 times the size of ChEX. Notably, none of the baseline models is capable of performing all the tasks covered by ChEX.
We conduct a thorough analysis of ChEX’s interactive capabilities, demonstrating its proficiency in responding to specific user prompts.
2 Related Work
2.0.1 Medical Image-Text Models
One prevalent category of medical image-text models adopts a CLIP-style [58] framework, leveraging contrastive image-text learning [24, 89, 50, 4, 70, 38, 64, 72, 77]. While most works focus on global image-text alignment, only some works consider more localized elements such as individual sentences or image patches [24, 50, 4, 38, 64, 72].
Report generation has recently received significant attention [29, 21, 68, 69, 15, 56, 71, 82, 36, 26, 76, 55, 49, 73]. Some of these approaches target specific aspects such as pathologies [29, 21, 68], anatomical regions [69, 15], or both [36]. However, support for interactive user queries remains rare among report generation models, with only the model RGRG [69] enabling queries via bounding boxes and a few others [56, 71, 82] supporting textual queries, albeit without the ability to jointly predict bounding boxes and textual descriptions. Notably, the model OmniFM-DR [82] supports the prediction of bounding boxes for textual prompts but without describing the content therein.
The ability of responding to textual prompts characterizes visual question answering (VQA) models, with some multitask models [71, 82, 34, 83] supporting zero-shot medical VQA, while others require fine-tuning [88, 18, 67, 40, 12]. However, unlike ChEX, these models lack localization capabilities for their responses. Moreover, it is important to note that while these methods rely on question-answer pairs for training, ChEX indirectly acquires its interactive capabilities through multitask training. Although localization tasks have been integrated into multitask training for natural images [74, 87, 84], such approaches remain scarce in the medical domain, with OmniFM-DR [82] being a notable exception.
2.0.2 Prompt-based Localization
DETR [6] pioneered the use of token vectors for object detection, sparking a series of subsequent models following this approach [91, 90, 41, 47, 86]. However, these models typically employ tokens that are not input-dependent, often relying on learned or position-based tokens [41, 86, 47]. Visual grounding models [9, 8, 11, 37] predict bounding boxes for given textual phrases, relating them to image regions, and have been applied in medical imaging [7, 27]. GLIP [35, 87] unifies object detection and visual grounding using textual prompts, with applications emerging in the medical domain [17, 81]. Open-vocabulary object detection extends object detection models to unseen classes using textual prompts [80, 16, 85, 42, 62, 46]. Similarly, prompt-based segmentation techniques have been explored [33] with applications in the medical domain [25, 45].
3 ChEX: A Localized and Interactive Chest X-ray Description Model
Our model ChEX supports a wide range of tasks, spanning from localization to text generation, while considering different aspects like anatomical regions and pathologies. To facilitate efficient training and zero-shot inference across all tasks, ChEX employs a simple yet versatile architecture, outlined in Figure 2. First, an image encoder is used to extract patch tokens of the given chest X-ray, while each textual prompt is encoded using the prompt encoder, a frozen text encoder. Next, the prompt detector, a DETR[6]-style object detector, localizes the prompts within the image, predicting a set of bounding boxes for each prompt along with a single Region-of-interest (ROI) token vector per prompt. Finally, the sentence generator is conditioned on each ROI token independently as well as on all patches to predict a concise description of each ROI. Further details are given in the following Secs. 3.1, 3.2 and 3.3. For reproducibility, we provide comprehensive implementation details in Appendix 0.D.
3.1 Model Architecture
3.1.1 Image and Prompt Encoder
We use CLIP[58] with ViT-B/32[10] and the default text encoder, both pre-trained[70] on chest X-ray/report pairs from MIMIC-CXR[30]. Instead of using the CLS token, we use all patch tokens from the ViT and project them individually into the shared image-text space using the pre-trained projector from CLIP. Given textual prompts, each of them is embedded independently using the prompt encoder, i.e. the text encoder from CLIP, and projected to the shared image-text space. During training, we freeze the complete prompt encoder and the image encoder up to the 8th encoder layer.
3.1.2 Prompt Detector
Using the prompt tokens from the output of the prompt encoder and the patch tokens from the image encoder, the prompt detector independently localizes each prompt by predicting bounding boxes per prompt, following the DETR [6] decoder with 6 layers. To support multiple boxes per prompt, we adopt [85] and additively combine each prompt token with learned tokens, leading to a total of decoder query tokens. After decoding, we apply an MLP-based box predictor on each token feature to predict individual bounding boxes and box scores. We then use the bounding box parameters as a regional bias and compute box features by applying Gaussian ROI pooling [52] on the patch tokens. Additionally, we introduce a random skip connection to provide a direct path from the decoder layers. Finally, we compute the ROI tokens by aggregating the box features per prompt using a weighted average based on box scores. When using bounding box queries instead of textual prompts, ROI features are directly computed with Gaussian ROI pooling on patch tokens using the given bounding box parameters and bypassing all decoder layers.
3.1.3 Sentence Generator
For sentence generation, we use the GPT2-medium [59] model pre-trained on PubMed abstracts and condition it on each ROI token independently using P-tuning v2 [43] with MLP projection and without freezing parameters. To incorporate additional global context, we apply a post decoder comprising three transformer decoder layers, with the ROI tokens as queries and the patches as keys and values before feeding the features into the GPT2 model.
3.2 Multitask Training
We train ChEX in a multitask setting where each sample may provide targets for different aspects of the image. To enable training in such a setting, we use three different types of queries depending on the available targets per sample: (i) pathology tokens, (ii) anatomy tokens, and (iii) sentence tokens.
Pathology tokens are employed when pathology class labels or bounding boxes are available with a sample. We encode the textual prompt of each pathology, which then serves as a prompt token for the prompt detector. The generated ROI tokens are trained via a (supervised) contrastive loss, incorporating both positive and negative prompts for each pathology. In positive pairs, the ROI token for a pathology is paired with its positive or negative pathology prompt, depending on the presence or absence of this pathology. Conversely, negative pairs include the opposing prompt alongside all positive and negative prompts of all other pathologies. If target bounding boxes are available, we use a modified version of the DETR loss [6]. We iteratively apply Hungarian matching per pathology until all predicted boxes are matched and then apply the L1 and gIoU losses from the original DETR loss. Box scores are trained by applying the focal loss [39] with positive targets for boxes matched in the first iteration.
Anatomy tokens are employed when anatomical region targets are available, including bounding boxes, associated pathology labels, and report sentences per region. We use the anatomical region names as textual prompts, training their bounding boxes using L1 and gIoU losses, while box scores remain untrained. If multiple target boxes are available per region, we apply Hungarian matching. For regions with a single bounding box, we match all its predicted boxes with the target box. Similar to the approach adopted for pathological tokens, we apply a supervised contrastive loss using positive and negative pathology prompts. For positive pairs, we align the ROI token of a region with the set of all positive pathology prompts of pathologies present in this region, along with negative prompts of pathologies not present. All other pathology prompts are used as negatives. Additionally, we train the sentence generator conditioned on anatomy features (i.e. ROI tokens) using the target sentences assigned to each anatomical region. We apply the standard autoregressive language modeling loss considering only regions where target sentences are available. If multiple target sentences are available, we concatenate them. For each sample, we randomly decide whether the anatomy features (i.e. ROI tokens) are computed using textual anatomy prompts or queries based on target bounding boxes of anatomical regions.
Sentence tokens are employed when report sentences are available. In such cases, the encoded sentences serve as prompt tokens. Since target bounding boxes and pathology labels are absent for sentences, we rely solely on the sentences themselves. We use a contrastive loss, where the extracted ROI token for each sentence and the encoded sentence itself form positive pairs, while other encoded sentences from the same and different samples serve as negatives. Additionally, we apply the standard CLIP loss [58] on average-pooled patch and sentence features. Furthermore, we train the sentence generator to reconstruct the original sentence based on its ROI token.
3.3 Zero-shot Inference
During inference, textual prompts like predefined pathology names, sentences, or user queries are encoded, and the prompt detector predicts bounding boxes and ROI tokens. These tokens are used by the sentence generator to predict descriptions of the detected regions. If regions require classification, such as for object detection or region classification, ROI tokens are classified based on cosine similarity with encoded prompts. When bounding boxes are given as queries, they’re used for Gaussian ROI pooling, omitting other parts of the prompt detector and encoder if no textual prompts are provided. For full report generation, we use pre-defined sets of textual prompts and concatenate the predicted descriptions.
4 Experimental Setup and Evaluation
4.1 Training Dataset and Pre-processing
We train on the frontal chest X-rays from MIMIC-CXR[30, 31, 14] and VinDr-CXR[54, 53, 14]. MIMIC-CXR comes paired with radiology reports, of which we use the findings and impression sections and split them into individual sentences. We use additional supervision for the MIMIC-CXR images provided by the Chest ImaGenome (CIG) [78, 79, 14] dataset. It includes bounding boxes for 29 unique anatomical regions and additionally assigns report sentences as well as 53 unique findings and pathology labels to each of these regions. VinDr-CXR includes bounding boxes for 22 unique findings and pathologies. Overall, we train on 227,382 images from MIMIC-CXR and 15,000 images from VinDr-CXR, from their official train splits, but oversample VinDr-CXR samples to simulate equal size of both datasets. We randomly crop and resize all images to a resolution of and then apply random horizontal flips, random affine transformations, contrast/brightness jittering, and random Gaussian blurring.
4.2 Benchmark Tasks
We evaluate the zero-shot performance of our model ChEX across the 9 chest X-ray tasks shown in Tab. 1. For details on the evaluation and dataset preparation, we refer to Appendix 0.E. While no task-specific fine-tuning is applied, post-processing (e.g. box suppression and scaling of boxes) for ChEX and all baselines is adjusted to consider differences in annotation practices of datasets.
Task Dataset | #Samples | #Classes | Evaluation Metrics |
Sentence Grounding (SG): Predicting bounding boxes for given sentences | |||
MS-CXR[4] | 169 | none | mIoU, mAP |
Pathology Detection (OD): Object detection of pathologies | |||
VinDrCXR[54] | 1,500 | top 15 | mAP |
NIH ChestXray (NIH8)[75] | 448 | 8 | mAP |
MS-CXR[4] | 169 | 8 | mAP |
Region Classification (RC): Classifying regions defined by given bounding boxes | |||
MS-CXR[4] | 169 | 8 | AUROC |
Chest ImaGenome (CIG)[78] | 3,402 | 53 | weighted AUROC (wAUROC) |
Region Explanation (RE): Predicting descriptions for regions defined by bounding boxes | |||
MS-CXR[4] | 169 | none | METEOR[1] Mic-F1-14†, Mac-F1-14† |
Chest ImaGenome (CIG)[78] | 3,402 | none | METEOR[1] Mic-F1-14†, Mac-F1-14† |
Full Report Generation (RG): Predicting full reports from chest X-rays | |||
MIMIC-CXR[30] | 3,082 | none | METEOR[1] Mic-F1-14†, Mac-F1-14†, Ex-F1-14† Mic-F1-5+†, Mac-F1-5+† |
4.3 Benchmark Baselines
We consider a wide variety of specialist and generalist baselines in our benchmark. For SG, we consider the standard supervised visual grounding (SupVG) model TransVG[9] and the generative multitask model OmniFM-DR[82]. For OD and RC tasks, we compare against standard supervised object detection (SupOD) models including Faster R-CNN[63] and DETR-style models [6, 47, 91, 90], and weakly-supervised object detection (WSupOD) including ADPD[51] and CheXNet[60], each of them with different (pre-trained) backbones. Additionally, we study the zero-shot capabilities of standard CLIP-style models for chest X-rays (BioVIL[4] and CheXzero[70]) on all SG, OD, and RC tasks. To the best of our knowledge, there is only one model supporting the RE tasks out-of-the-box, namely RGRG[69]. We also use the two CLIP-style baselines as image-text retrieval models and evaluate their performance on these tasks. For RG, we compare against several report generation models, including recent models like MAIRA-1[26], Med-PaLM M [71], Prompt-MRG[29], and RaDialog[56].
5 Results
We compare ChEX with SOTA models (Sec. 5.1). Additionally, we study how ChEX reacts to (interactive) user queries (Secs. 5.2 and 5.3) and show its interpretable and customizable report generation capabilities (Sec. 5.4). ChEX is competitive with SOTA models while providing a high degree of interpretability and interactivity, therefore offering a promising path towards clinical application.
5.1 Comparison with SOTA
Fig. 3 and Tab. 2 provide an overview of the performance of our model ChEX compared to the best baselines, including specialized SOTA and common multitask models. On 8 of the 9 tasks, our model ChEX is competitive (within 1-std) or better than the best baseline on at least one metric. Only on pathology detection (OD) on VinDR-CXR is it outperformed by a specialized supervised object detection model. Note that none of the baselines is capable of performing all the tasks, as most of them are either focused solely on localization or generative tasks, but not both. Only the contrastive image-text (i.e. CLIP-style) models can perform a wide range of tasks but rely on retrieval for the generative tasks, thus showing poor performance on these tasks. Overall, our model ChEX shows excellent performance on a wide range of tasks, covering both localization as well as text generation, and is capable of replacing specialized models without major performance drops on most tasks. Detailed results, ablation studies, and further analysis can be found in the Appendices 0.A, 0.B and 0.C.
Task | Metric | ChEX | SupOD | WSupOD | SupVG | Contrastive | Generative | ||||||
Sentence Grounding (SG) | |||||||||||||
MS-CXR | [mIoU] | 1.45 | – | – | 1.53 | 1.31 | |||||||
[mAP] | 2.21 | – | – | 2.63 | 1.37 | – | |||||||
Pathology Detection (OD) | |||||||||||||
VinDrCXR | [mAP] | 0.95 | 1.20 | 0.88 | – | 0.25 | – | ||||||
NIH8 | [mAP] | 1.05 | 0.82 | 0.88 | – | 0.26 | – | ||||||
MS-CXR | [mAP] | 1.38 | 1.42 | 1.06 | – | 0.52 | – | ||||||
Region Classification (RC) | |||||||||||||
MS-CXR | [AUROC] | 2.80 | 2.55 | 3.41 | – | 2.80 | – | ||||||
CIG | [wAUROC] | 0.36 | 0.22 | 0.21 | – | 0.32 | – | ||||||
Region Explanation (RE) | |||||||||||||
MS-CXR | [Mic-F1-14] | 2.24 | – | – | – | 1.41 | 2.50 | ||||||
[Mac-F1-14] | 1.54 | – | – | – | 0.67 | 2.00 | |||||||
[METEOR] | 0.54 | – | – | – | 0.36 | 0.78 | |||||||
CIG | [Mic-F1-14] | 0.43 | – | – | – | 0.38 | 0.44 | ||||||
[Mac-F1-14] | 0.35 | – | – | – | 0.15 | 0.19 | |||||||
[METEOR] | 0.13 | – | – | – | 0.03 | 0.10 | |||||||
Full Report Generation (RG) | |||||||||||||
MIMIC-CXR† | [Mic-F1-14] | 0.51 | – | – | – | – | |||||||
[Mac-F1-14] | 0.51 | – | – | – | – | ||||||||
[Ex-F1-14] | 0.42 | – | – | – | – | ||||||||
[Mic-F1-5+] | 0.56 | – | – | – | – | ||||||||
[Mac-F1-5+] | 0.57 | – | – | – | – | ||||||||
[METEOR] | 0.10 | – | – | – | – |
5.1.1 Localization and Region Classification Tasks
In sentence grounding (SG), ChEX demonstrates similar performance to the only generative model OmniFM-DR[82] and the best SupVG model TransVG[9] (with CheXzero[70] backbone), where TransVG shows an advantage on the mIoU metric. Notably, TransVG was trained explicitly on this task. ChEX outperforms TransVG without the CheXzero backbone by mAP and is within 1-std on mIoU. The superiority of ChEX over contrastive models underscores the importance of bounding box supervision for achieving accurate localization.
ChEX is competitive on 2 of 3 pathology detection (OD) tasks, namely NIH8 and MS-CXR. On VinDr-CXR, where bounding box supervision is available, the SupOD model Faster R-CNN [63], trained on this dataset, performs best. o-DETR[91] performs slightly worse but still outperforms ChEX, while other DETR-style models are on-par or inferior to ChEX. ChEX doubles the performance of the best WSupOD model Loc-ADPD[51] (with BioVIL backbone). On NIH8, all SupOD models perform very poor (ChEX almost doubles their best performance), while the WSupOD model Loc-ADPD is competitive with ChEX. On MS-CXR, the best SupOD baseline o-DETR (with BioVIL backbone), trained on VinDr-CXR, and the best WSupOD model Loc-ADPD are both within 1-std of ChEX. Zero-shot contrastive models perform poorly, as they rely on thresholding on noisy similarity maps.
On region classification (RC) tasks, ChEX clearly outperforms all baselines. The best baseline on MS-CXR is the Faster R-CNN with BioVIL backbone, trained on VinDr-CXR, and is outperformed by , while on CIG the best baseline is the contrastive model BioVIL, which is outperformed by .
5.1.2 Generative Tasks
Region explanation (RE) can only be performed by a single generative model, namely RGRG[69]. Other report generation models can only predict the full report but cannot provide descriptions for specified regions. On MS-CXR, our model ChEX performs similar to RGRG on Mic-F1-14 and METEOR but improves by on Mac-F1-14. On CIG, ChEX outperforms RGRG by on Mic-F1-14, by on Mac-F1-14, and by on METEOR. This result is particularly noteworthy considering RGRG’s explicit training for this task and highlights the importance of the contrastive losses in ChEX. Sentence-retrieval with contrastive baselines performs poorly on these tasks.
For full report generation (RG) on MIMIC-CXR, ChEX sets a new state-of-the-art on the Ex-F1-14, Mic-F1-5+, and Mac-F1-5+ metrics. Considering the commonly used metric Mic-F1-14, the SOTA model MAIRA-1[26] – having 7B parameters – outperforms ChEX – having only 1B parameters – by only . More notably, ChEX outperforms the more than ten times larger 12B parameter Med-PaLM M by and is only outperformed by by the 80 times larger 84B parameter version. Limitations of ChEX can be observed on the Mac-F1-14 metric, where ChEX is outperformed by the SOTA model Med-PaLM M 84B by but also by smaller models like Prompt-MRG[29] (by ) and RaDialog-RG[56] (by ). ChEX, however, improves over RGRG by , indicating that this is a limitation of training based on anatomical regions and is caused by the poor performance of some rarer pathologies (we refer to Appendix 0.C). On the Ex-F1-14 metric, ChEX outperforms the best baseline Prompt-MRG by . Most larger models do however not report this metric. On the Mic-F1-5+ and Mac-F1-5+ metrics, ChEX outperforms the large SOTA model MAIRA-1 by and , respectively. In language-based metrics such as METEOR, ChEX shows relatively low performance ( lower than the best and lower than the second-best model) due to not being explicitly trained for full report generation but generating reports based on a set of queries. While this approach offers advantages (see Sec. 5.4), it may result in reports differing from reference reports. However, it’s important to note that language-based metrics offer limited insights into the factual correctness of generated reports [49]. Overall, ChEX showcases strong report generation performance, achieving results that are close or even better than SOTA models of up to 80 times the size of ChEX.
5.2 Interactive Prompting
[…]
5.3 Improvement through Precise Interactive Prompting
[…]
5.4 Interpretable and Customizable Report Generation
ChEX supports automatic full report generation using a pre-defined set of textual prompts based on which the model localizes and describes relevant regions. Unlike typical report generation models, the reliance on prompt sets for report generation offers high flexibility as the prompt sets can be customized without the need for re-training. We study (cf. Appendix 0.A) the effect of using only pathologies, anatomical regions, or both as prompt sets. The choice of the prompt set enables balancing precision and recall of the model. While all studied prompt sets lead to results competitive with the baselines, with Mic-F1-14 ranging from to , using both prompt sets leads to optimal performance.
In Fig. 7, we show an example of a generated report. The prediction of bounding boxes for predicted sentences enables a high degree of transparency and interpretability – not provided by most report generation models – thus simplifying correctness checking and enabling an optimal clinical workflow.
6 Discussion
6.0.1 Importance of Semantically Meaningful Visual Features
Many recent works on radiology report generation heavily focus on the utilization of large language models (LLMs) [26, 71, 82, 56]. We demonstrate that even smaller language models can achieve competitive performance when prioritizing improvements in image understanding aspects. Specifically, our approach places emphasis on semantic regions, leveraging bounding box supervision to enhance prediction quality. We argue that a synergistic combination of optimal image encoding with regional focus along with the deductive powers and knowledge of LLMs presents a promising path for improving report generation. Furthermore, our proposed training strategy enables the integration of multiple tasks and datasets, offering a viable approach to realizing this path.
6.0.2 Interpretability and Interactivity for Clinical Practice
While the performance of report generation models is improving, even occasional inaccuracies in their predictions can limit their clinical applicability. We argue that a high degree of interpretability and the integration of medical experts in the generation process, i.e. interactivity, offers a more promising path to clinical application than solely improving prediction quality. Our model ChEX showcases a unique combination of interpretability and interactivity – not provided by other models – by providing bounding boxes for generated descriptions while enabling user guidance through textual prompts and bounding box queries.
6.0.3 Limitations
While our training approach enables effective utilization of the available datasets, it also leads to some limitations. The used datasets do not provide question-answer pairs, so textual queries are either based on pre-defined prompts or report sentences, and answers are always based on report sentences. This limits the types of supported textual queries, mainly regional or pathology hints, or both. At the same time, answers are plain descriptions of the queried regions or pathologies, while specific answers to more complicated questions (e.g. comparing regions) are not supported. Also, using report sentences as answers can lead to the hallucination of comparisons with previous images, although only a single image is used, a phenomenon common to report generation models [26, 2, 61]. Future work may use instruction tuning to tackle these issues.
6.0.4 Conclusion
We proposed ChEX, a model for predicting visually grounded textual descriptions of chest X-rays based on user queries. Our analysis underscores ChEX’s competitive performance against SOTA models across 9 tasks, and its responsiveness to user prompts, therefore laying a foundation for future advancements in interactive and localized text generation models.
Appendix 0.A Ablation Studies and Further Analysis
0.A.1 Multitask Training: Token and Target Types
ChEX | Token Types | Target Types | |||||||||
P-only | A-only | S-only | No-P | No-A | No-S | No box | No cls | No sent | |||
Pathology boxesV | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
Pathology classesV | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
Anat. boxesM | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
Anat. patho classesM | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
Anat. sentencesM | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
SentencesM | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
Sentence Grounding (SG) | |||||||||||
MS-CXRM | [mIoU] | 47.52 | 42.34 | 45.07 | 9.00 | 23.90 | 34.92 | 52.97 | 3.68 | 44.38 | 52.97 |
[mAP] | 44.47 | 37.92 | 40.47 | 0.94 | 18.89 | 19.55 | 50.66 | 3.71 | 36.43 | 50.11 | |
Pathology Detection (OD) | |||||||||||
VinDrV | [mAP] | 14.12 | 13.25 | 4.60 | 0.06 | 1.16 | 14.15 | 14.38 | 0.22 | 10.35 | 14.38 |
NIH8N | [mAP] | 11.14 | 7.02 | 4.40 | 0.22 | 2.23 | 5.38 | 7.96 | 0.56 | 6.70 | 7.96 |
MS-CXRM | [mAP] | 16.60 | 13.52 | 10.90 | 0.26 | 8.83 | 5.51 | 14.11 | 0.52 | 14.96 | 14.11 |
Region Classification (RC) | |||||||||||
MS-CXRM | [AUROC] | 82.33 | 74.02 | 79.86 | 49.49 | 76.58 | 69.10 | 80.89 | 46.87 | 79.89 | 75.67 |
CIGM | [wAUROC] | 70.46 | 59.36 | 69.61 | 50.08 | 68.58 | 62.28 | 69.61 | 58.03 | 70.16 | 69.89 |
Region Explanation (RE) | |||||||||||
MS-CXRM | [Mic-F1-14] | 49.97 | 45.44† | 39.47 | 10.94 | 22.49 | 35.37 | 42.72 | 5.53 | 35.02 | 43.16† |
[Mac-F1-14] | 20.50 | 17.99† | 20.29 | 1.48 | 15.19 | 9.37 | 20.01 | 2.61 | 15.17 | 14.40† | |
[METEOR] | 8.79 | 6.91† | 9.02 | 0.00 | 6.13 | 5.69 | 7.74 | 2.96 | 7.77 | 5.73† | |
CIGM | [Mic-F1-14] | 53.34 | 35.26† | 51.25 | 10.96 | 47.51 | 28.76 | 51.84 | 23.47 | 53.22 | 48.93† |
[Mac-F1-14] | 29.13 | 16.00† | 27.65 | 1.66 | 24.14 | 14.82 | 28.20 | 10.73 | 29.18 | 24.36† | |
[METEOR] | 10.18 | 5.48† | 9.57 | 0.00 | 7.95 | 4.70 | 9.79 | 3.47 | 10.08 | 8.89† | |
Full report generation (RG) | |||||||||||
MIMIC-CXRM | [Mic-F1-14] | 52.32 | 50.10† | 49.64 | 16.73 | 43.44 | 46.64 | 49.91 | 40.68 | 51.69 | 51.18† |
[Mac-F1-14] | 32.56 | 26.76† | 29.87 | 3.22 | 27.60 | 25.35 | 31.32 | 22.83 | 30.75 | 29.81† | |
[Ex-F1-14] | 58.76 | 59.04† | 57.60 | 50.25 | 50.10 | 55.02 | 57.39 | 55.93 | 57.39 | 58.44† | |
[Mic-F1-5+] | 61.03 | 58.16† | 57.09 | 44.28 | 52.10 | 54.43 | 57.77 | 49.79 | 60.58 | 58.63† | |
[Mac-F1-5+] | 55.85 | 48.65† | 52.31 | 22.81 | 49.96 | 50.20 | 52.55 | 40.75 | 54.54 | 51.62† | |
[METEOR] | 13.26 | 11.13† | 12.45 | 0.00 | 12.69 | 10.97 | 13.05 | 8.01 | 13.55 | 11.99† |
† We additionally trained the sentence generator afterwards (with anatomy and report sentences).
0.A.1.1 Token Types
We study the effect of different forms of supervision (i.e. token and target types) during multitask training (Tab. 3). In the default case, ChEX is trained with all token types (pathology, anatomy, sentence) and all target types (bounding boxes, pathology classification, sentences). Using only pathology tokens (with bounding box and class targets), i.e. only samples from VinDr-CXR, leads to performance reductions on SG, OD, and RC tasks. When additionally using sentence supervision for training only the sentence generator, the performance on the RE task on MS-CXR and the RG task is only slightly degraded compared to ChEX, while there is a huge performance drop on the RE task on CIG. Using only anatomy tokens (with bounding boxes, pathology labels, and sentences) leads to a further performance drop on OD, RE on MS-CXR, and RG, while performance on SG and RE on CIG is improved. Using only sentence tokens leads to very poor performance on all tasks. Similarly, using no pathology tokens or no anatomy tokens leads to very poor performance on most tasks, with the exception that using no anatomy tokens leads to good performance on VinDR-CXR OD. Using no sentence tokens only leads to small performance drops on most tasks and improves performance on SG as well as on VinDr-CXR OD, while only leading to meaningfully performance drops on the other two OD tasks.
Overall, this shows the importance of using all types of tokens. Having both pathology and anatomy tokens is especially relevant for good SG and OD performance, i.e. for accurate localization. Anatomy tokens are especially relevant for RC and RE tasks, due to the available of pathology labels and sentences associated with bounding boxes. Sentences tokens are the least relevant tokens and can degrade the performance on SG, i.e. they can lead to less accurate localization. However, they are relevant to achieve the best possible RE and RG performance, i.e. for optimal text generation, especially on the MS-CXR dataset and can also improve the OD performance on some datasets.
0.A.1.2 Target Types
We also study the use of all token types while excluding specific target types. As expected, excluding bounding box targets leads to very poor performance on all tasks involving localization, i.e. all SG and OD tasks. Huge performance drops can also be observed on RE and RG tasks, while drops are smaller on RC tasks. Excluding pathology classification labels (for pathology and anatomy tokens) reduces the performance mainly on OD tasks and RE on MS-CXR, while smaller performance drops can be observed on SG and RG. Almost no performance drops are observed on the RC tasks and RE on CIG. Excluding sentence targets (i.e. excluding sentence targets for anatomy tokens and all sentence tokens) leads to performance drops on most tasks, except on SG and OD on VinDr-CXR, where performance improves.
Overall, we identified the huge importance of bounding box targets for all tasks. Pathology labels are relevant mainly for OD tasks. Again, sentence supervision is relevant for sentence generation while being less relevant for other tasks. Contrastive sentence improves some OD tasks as well as MS-CXR-based RC and RE tasks, i.e. it helps with the understanding of pathology regions.
0.A.2 Prompt Sets for Report Generation
Prompt sets | MIMIC-CXR | ||||||
Anatomy | Pathology | [Mic-F1-14] | [Mac-F1-14] | [Ex-F1-14] | [Mic-P-14] | [Mic-R-14] | [METEOR] |
yes | yes | 51.04 | 30.61 | 57.05 | 42.99 | 62.80 | 13.90 |
yes | positive | 50.13 | 30.67 | 56.91 | 42.46 | 61.22 | 13.60 |
yes | no | 50.08 | 30.07 | 56.97 | 42.66 | 60.64 | 12.67 |
positive | yes | 52.37 | 32.69 | 58.78 | 45.09 | 62.44 | 13.28 |
positive | positive | 52.02 | 32.88 | 58.36 | 44.64 | 62.33 | 12.24 |
positive | no | 51.27 | 32.18 | 58.82 | 45.16 | 59.31 | 10.77 |
no | yes | 50.35 | 25.82 | 59.35 | 47.88 | 53.10 | 9.82 |
no | positive | 50.39 | 26.69 | 60.14 | 49.38 | 51.44 | 8.26 |
ChEX uses pre-defined prompt sets to enable full report generation (RG). In Tab. 4, we study the effect of using pathology and anatomical region prompts. We additionally study the effect of excluding regions without pathologies (based on contrastive classification of the CheXpert[28] pathologies). Overall, all studied prompt sets lead to results competitive with the baselines, with Mic-F1-14 ranging from to , but the selection of the prompt set influences the relation between precision (Mic-P-14) and recall (Mic-R-14). Using only pathology prompts leads to a high precision but achieves much smaller macro F1 scores than also using anatomy prompts (with positive filtering). This is explained by the poor performance on a few pathologies (adding an anatomical prompt with positive filtering to the prompt set achieves -fold improvement for enlarged cardiomediastinum and -fold improvement for lung lesions). Therefore, using anatomical prompts can spot additional pathologies that may be missed by pathology prompts. Similarly, using only anatomy prompts leads to a small recall because pathologies may again be missed. Using both prompt sets leads to the best results. However, considering all anatomical regions can lead to overprediction (lower precision) and, therefore, lower overall F1 scores. We found that using positive filtering with anatomy prompts (i.e. considering only anatomical regions with positive pathologies for generation) improves performance. The best setting for Mic-F1-14 uses positive filtering for anatomy prompts and no filtering for pathology prompts. We consider this the default setting of ChEX, and all reported results from the main paper follow this setting. The best Mac-F1-14 can be achieved using positive filtering for both prompt sets. The best METEOR score and recall can be achieved using no positive filtering with both prompt sets, while the highest precision is achieved using only pathology prompts with positive filtering. We highlight that changing the prompt set does not require re-training.
0.A.3 Learned Regional Bias for Pathology Prompts
In Fig. 8, we study the regional bias of pathology prompts in negative images. Cardiomegaly prompts typically overlap with the cardiac silhouette, while most other pathologies mostly overlap with lung regions, as expected. More noteworthy, pleural effusion is typically searched in the lower lung zones, while pneumothorax is searched in the upper lung zones, which aligns well with the typical locations of these pathologies in chest X-rays. This indicates that our model has learned a meaningful regional bias for pathologies.
0.A.4 Query Type for Region Classification and Explanation
Chest ImaGenome (CIG) | ||
Region Classification (RC) | Region Explanation (RE) | |
Query type | [wAUROC] | [Mic-F1-14] |
Boxes (default task) | 70.460.36 | 53.340.43 |
Textual prompt | 70.760.37 | 54.700.41 |
Tab. 5 shows the effect of using bounding boxes versus textual prompts as queries for RC and RE on the CIG dataset, i.e. for anatomical regions. When using textual prompts, i.e., the names of anatomical regions, instead of their bounding boxes (the default setting), the performance of ChEX slightly improves. We assume that these improvements are caused by the additional context provided by the region prompts. Overall, the differences are marginal, showing that our model ChEX is robust to the choice of user query.
0.A.5 Sensitivity to Prompt Modifications
[…]
Appendix 0.B Qualitative Examples and Failure Cases
Appendix 0.C Detailed Benchmark Results
0.C.1 Pathology Detection (OD) Tasks
Pathology Detection (OD) | |||
VinDrV | NIH8N | MS-CXRM | |
[mAP] | [mAP] | [mAP] | |
ChEXVM | 14.120.95 | 11.141.05 | 16.601.38 |
Supervised Object Detection | |||
FRCNN[63] V | 18.211.20 | 6.690.82 | 15.131.22 |
FRCNN[63] (BioVIL[4]) VM | 16.171.31 | 6.560.90 | 15.641.16 |
FRCNN[63] (CheXzero[70]) VM | no convergence | ||
DETR-R50[6] V | 11.910.96 | 5.900.85 | 8.940.35 |
DETR-BioVIL[6, 4] VM | 3.820.59 | 2.460.46 | 1.940.48 |
DETR-CheXzero[6, 70] VM | 1.260.39 | 1.180.34 | 0.180.15 |
DETR-DC5-R50[6] V | 11.940.89 | 4.210.61 | 10.300.87 |
DETR-DC5-BioVIL[6, 4] VM | 13.401.43 | 4.480.61 | 8.120.69 |
Conditional DETR-DC5-R50[47] V | 13.350.82 | 3.930.57 | 13.501.09 |
Conditional DETR-DC5-BioVIL[47, 4] VM | 14.861.33 | 4.970.67 | 11.990.98 |
Deformable DETR-R50[90] V | 11.491.13 | 2.940.50 | 10.650.88 |
Deformable DETR-BioVIL[90, 4] VM | 12.270.94 | 4.890.63 | 10.260.82 |
o-Deformable-DETR-R50[91] V | 16.391.13 | 4.260.63 | 14.861.20 |
o-Deformable-DETR-BioVIL[91, 4] VM | 15.971.00 | 5.380.75 | 15.831.42 |
Weakly-Supervised Object Detection | |||
CheXnet[60] N | 1.830.25 | 7.120.74 | 7.470.60 |
CheXnet[60] (BioVIL[4]) NM | 1.620.17 | 4.930.41 | 4.660.51 |
CheXnet[60] (CheXzero[70]) NM | 0.300.04 | 0.680.18 | 1.400.53 |
MIL-ADPD[51] M | 4.300.47 | 8.140.77 | 7.210.34 |
MIL-ADPD[51] (BioVIL[4]) M | 2.190.27 | 3.290.34 | 2.200.34 |
Loc-ADPD[51] M | 6.690.42 | 11.330.90 | 14.710.82 |
Loc-ADPD[51] (BioVil[4]) M | 7.440.41 | 11.890.88 | 16.561.06 |
Zero-Shot Contrastive | |||
BioVIL[4] M | 2.820.25 | 2.630.26 | 7.150.52 |
CheXzero[70] M | 0.540.06 | 2.390.24 | 2.810.25 |
0.C.2 Sentence Grounding (SG) Task
Sentence Grounding (SG) | ||
MS-CXRM | ||
[mIoU] | [mAP] | |
ChEXVM | 47.521.45 | 44.472.21 |
Zeroshot Contrastive Models | ||
BioVIL[4] (masks) M | 22.751.26 | – |
BioVIL[4] (boxes) M | 28.571.31 | 18.621.37 |
CheXzero[70] (masks) M | 11.940.59 | – |
CheXzero[70] (boxes) M | 15.450.67 | 5.940.64 |
Multitask Generative Models | ||
OmniFM-DR[82] VMO | 46.2 | – |
Supervised Visual Grounding | ||
TransVG[9] M | 48.811.45 | 37.652.61 |
TransVG[9] (BioVIL[4]) M | 52.131.73 | 41.242.73 |
TransVG[9] (CheXzero[70]) M | 53.511.53 | 44.052.63 |
0.C.3 Region Classification (RC) Tasks
Region Classification (RC) | ||
MS-CXRM | Chest ImaGenomeM | |
[AUROC] | [wAUROC] | |
ChEX VM | 82.332.80 | 70.460.36 |
Supervised Image Classifiers (ROI-pooled) | ||
CheXnet[60] N | 61.463.41 | 59.030.21 |
CheXnet[60] (BioVIL[4]) NM | 55.193.29 | 58.120.21 |
CheXnet[60] (CheXzero[70]) NM | 57.542.91 | 60.020.21 |
ROI-Pool Object Detectors (Classification Heads) | ||
Faster R-CNN[63] V | 75.662.64 | 58.280.22 |
Faster R-CNN[63] (BioVIL[4]) VM | 76.132.55 | 56.640.19 |
Zeroshot Contrastive Models | ||
BioVIL[4] ROI Pool M | 67.412.80 | 66.960.32 |
BioVIL[4] Cropped M | 59.563.09 | 60.410.26 |
ChexZero[70] ROI Pool M | 59.971.14 | 65.730.35 |
ChexZero[70] Cropped M | 60.533.45 | 61.800.28 |
0.C.4 Region Explanation (RE) Tasks
Region Explanation (RE) | |||||
MS-CXRM | |||||
[Mic-F1-14] | [Mac-F1-14] | [Mic-F1-5+] | [Mac-F1-5+] | [METEOR] | |
ChEX VM | 49.972.24 | 20.501.54 | 62.541.50 | 44.952.23 | 8.790.54 |
Generative Models | |||||
RGRG[69] M | 48.972.50 | 16.372.00 | 60.392.21 | 38.342.23 | 8.150.78 |
Zeroshot Contrastive Models (Nearest Neighbor, ROI pooled) | |||||
BioVIL[4] M | 5.861.41 | 3.690.67 | 12.792.25 | 10.951.93 | 3.820.03 |
CheXzero[70] M | 5.411.30 | 3.400.78 | 11.002.25 | 9.021.83 | 3.470.45 |
Region Explanation (RE) | |||||
Chest ImaGenomeM | |||||
[Mic-F1-14] | [Mac-F1-14] | [Mic-F1-5+] | [Mac-F1-5+] | [METEOR] | |
ChEX VM | 53.340.43 | 29.130.35 | 44.530.40 | 39.190.37 | 10.180.13 |
Generative Models | |||||
RGRG[69] M | 45.260.44 | 20.880.19 | 31.930.31 | 27.860.32 | 7.880.10† |
Zeroshot Contrastive Models (Nearest Neighbor, ROI pooled) | |||||
BioVIL[4] M | 24.400.38 | 7.150.10 | 9.420.22 | 7.450.16 | 3.820.03 |
CheXzero[70] M | 21.860.35 | 8.930.15 | 7.780.21 | 6.990.17 | 3.680.04 |
†Results differ to the RGRG paper [69], because RGRG uses a different test split and computes scores only for regions that where selected for prediction by the model.
0.C.5 Full Report Generation (RG) Task
Full Report Generation (RG) | |||||
MIMIC-CXRM | |||||
[Mic-F1-14] | [Ex-F1-14] | [Mac-F1-14] | [Mic-F1-5+] | [Mac-F1-5+] | |
ChEX VM | 52.320.51 | 58.760.42 | 32.560.51 | 61.030.56 | 55.850.57 |
RGRG[69] M | 49.8 | 44.7 | 54.7 | 46.44 | |
CvT-212DistilGPT2[55] M | 44.2 | 38.4 | 30.7 | – | |
M2 Trans w/NLL+BS+fCEN[49] M | – | – | – | 56.7 | – |
METransformer [76] M | – | 31.1 | – | – | – |
Med-PaLM M (12B) [71] MO | 51.41 | – | 37.31 | 56.54† | 50.57† |
Med-PaLM M (84B) [71] MO | 53.56 | – | 39.83 | 57.88† | 51.60† |
Med-PaLM M (562B) [71] MO | 51.60 | – | 37.81 | 56.28† | 49.86† |
MAIRA-1[26] M | 55.7 | – | 38.6 | 58.8 | 51.7 |
Prompt-MRG[29] M | – | 47.6 | 38.1 | – | – |
OmniFM-DR[82] MO | – | 33.3 | – | – | |
RaDialog-INS[56] MO | – | – | 38.6 | – | – |
RaDialog-RG[56] MO | – | – | 39.4 | – | – |
COOMG[15] | – | 34.5 | – | – | – |
ORGAN[21] | – | 38.5 | – | – | – |
† Limited comparability as F1-5 was reported instead of F1-5+, i.e. uncertain cases were treated as positive instead of negative.
* Reported number taken from the Prompt-MRG[29] paper.
Full Report Generation (RG) | |||||
MIMIC-CXRM | |||||
[BLEU-1] | [BLEU-4] | [METEOR] | [ROUGE-L] | [CIDEr] | |
ChEX VM | 29.580.24 | 6.520.11 | 13.260.10 | 19.490.10 | 18.490.79 |
RGRG[69] M | 37.3 | 12.6 | 16.8 | 26.4 | 49.5 |
CvT-212DistilGPT2[55] M | 39.2 | 12.4 | 15.3 | 28.5 | 36.1 |
M2 Trans w/ NLL+BS+fCEN[49] M | – | 11.4 | – | – | 50.9 |
METransformer[76] M | 38.6 | 12.4 | 15.2 | 29.1 | 36.2 |
Med-PaLM M (12B)[71] MO | 30.90 | 10.43 | – | 26.16 | 23.43 |
Med-PaLM M (84B)[71] MO | 32.31 | 11.31 | – | 27.29 | 26.17 |
Med-PaLM M (562B)[71] MO | 31.73 | 11.50 | – | 27.49 | 25.27 |
MAIRA-1[26] M | 39.2 | 14.2 | 33.3 | 28.9 | – |
Prompt-MRG[29] M | 39.8 | 11.2 | 15.7 | 26.8 | – |
OmniFM-DR[82] MO | – | 11.0 | 14.0 | 26.5 | – |
RaDialog-INS[56] MO | 34.0 | 9.7 | 13.6 | 27.0 | – |
RaDialog-RG[56] MO | 34.6 | 9.5 | 14.0 | 27.1 | – |
ITA[73] M | 39.5 | 12.1 | 14.7 | 28.4 | – |
COOMG[15] M | 34.6 | 10.4 | 13.7 | 27.9 | – |
COOMG-RL [15] M | 36.3 | 12.4 | 12.8 | 29.0 | – |
ORGAN[21] | 40.7 | 12.3 | 16.2 | 29.3 | – |
BioViL-T[2] M | – | 9.2 | – | 29.6 | – |
F1 | |||
Pathology/Finding | ChEX | RGRG[69] | MAIRA-1[26] |
Cardiomegaly | 66.621.08 | 62.4 | 64.0 |
Edema | 54.811.48 | 51.4 | 44.0 |
Consolidation | 15.272.10 | 7.8 | 20.0 |
Atelectasis | 45.991.43 | 54.6 | 41.3 |
Pleural Effusion | 69.111.04 | 56.0 | 68.9 |
Enlarged Cardiomediastinum | 6.811.74 | 0.3 | 11.9 |
Fracture | 0.000.00 | 0.0 | 24.9 |
Lung Lesion | 14.243.05 | 0.7 | 18.8 |
Lung Opacity | 52.321.06 | 26.8 | 49.8 |
Pleural Other | 0.000.00 | 0.2 | 14.7 |
Pneumonia | 19.301.55 | 16.2 | 18.3 |
Pneumothorax | 17.593.21 | 15.9 | 40.8 |
Support Devices | 69.400.92 | 70.9 | 84.5 |
No Finding | 24.362.97 | 63.2 | 38.6 |
Appendix 0.D Model Implementation Details
0.D.1 Model Architecture and Components
0.D.1.1 Prompt Detector
Fig. 12 provides an overview of the prompt detector. Given is an input of prompt tokens with . In order to enable the prediction of up to bounding boxes per prompt token , we utilize learned, randomly initialized, box tokens . We then additively condition these tokens on the queries resulting in a total of query tokens :
These tokens serve as the initial queries for the DETR decoder layers, while the patch tokens from the image encoder are used for keys and values. We use a total of 6 DETR-style transformer decoder layers. Self-attention is restricted, allowing attention only between tokens from the same query, i.e. with the same . The query tokens are only used as queries for the first layers, later layers use the output of the previous layer as queries. We additively combine the patch tokens with the corresponding positional encodings from the ViT before using them in the prompt detector. The outputs of the last decoder layer are used as preliminary ROI tokens . They are then used by the MLP-based box predictor to predict the corresponding bounding boxes, where we apply sigmoid to return relative image coordinates.
Next, we use Gaussian ROI pooling following [52]. For each bounding box, we define a 2D multivariate Gaussian with independent and components (i.e. with zero covariance) using the box center as mean and the box size as standard deviation. The probability density function (pdf) of this Gaussian is used in a weighted aggregation of the patch tokens (which have been projected by an MLP), resulting in a single feature vector per ROI . We randomly add the preliminary ROI tokens in of the cases but apply dropout to them before adding. Finally, these features are projected using an MLP with two hidden layers resulting in the ROI tokens . We use an additional MLP followed by sigmoid on these ROI tokens to predict box scores . Using these scores, we now aggregate the ROI tokens for each prompt to get the final (aggregated) ROI tokens :
0.D.1.2 Sentence Generator
Fig. 12 provides an overview of the sentence generator. It consists of two major components, the post decoder and the language model (GPT-2) with conditioning. The post decoder consists of three DETR decoder layers. For each ROI , we use the per-bounding-box ROI tokens and the aggregated ROI token as queries, leading to a total of queries per ROI. We randomly drop the per-bounding-box ROI tokens with a probability while always keeping the aggregated tokens . We use the patch tokens and the (projected) CLS token from the image encoder as keys and values. Self-attention is restricted, allowing attention only between queries for the same ROI . After the final decoder layer, we only keep the output features associated with the aggregated ROI token of each ROI , discarding the features of the per-bounding-box ROI tokens. The final output of the post decoder is computed by summing the outputs of the final decoder layer with the input features, introducing a skip connection. Before summing, we multiply the output features with a learned gating factor, which is initialized by zero.
For text generation, we condition the GPT-2 medium[59] language model on the output features of the post decoder. The descriptions of each ROI are predicted individually, treating each as an independent sample. We follow a similar approach as P-tuning v2 [43] but do not freeze any parameters of the language model. Specifically, we use an MLP to project the feature vector for each into a prefix for the language model. We use a prefix of five tokens and compute key and value features for each language model layer. During generation, we use a maximum length of 128 generated tokens.
0.D.1.3 Futher Implementation Details
For all components of the prompt detector and post decoder of the sentence generator, if not otherwise specified, we use the model dimension with hidden dimension for MLPs, dropout , attention dropout , drop-path[23] probability , and GELU[20] non-linearities. For attention layers we apply the layer norm before each operation and use layer scaling with initialization .
0.D.1.4 Number of Parameters
ChEX has an overall 1051M parameters of which we keep 123M frozen throughout our whole training. The ViT image encoder contains 88M, the text encoder 63M, the prompt detector 28M, and the sentence generator 872M parameters.
0.D.2 Supervision and Loss Functions
0.D.2.1 Pathology Tokens
[…]
Pathology Detection (Bounding Box Loss)
For the bounding boxes, we apply the modified DETR loss. The class match probability of a bounding box is computed as
where denotes the dot product and denotes the (non-aggregated) ROI token for each bounding box. Using the standard formulation of the box loss
we compute the matching cost for class , predicted box , and target box based on the box cost, class match probability of the box, and box score as
For each positive class , i.e. each class that has target bounding boxes, we iteratively apply Hungarian matching using the matching cost , where matching is applied individually for each . We start with an initial matching. If non-matched predicted boxes remain for that class, we apply matching again with the remaining predicted boxes and all target boxes until all predicted boxes are matched, allowing a target box to be matched with several predicted boxes. We apply the box loss to all pairs of matched predicted/target boxes. Additionally, we apply the focal loss to the box scores with positive targets for boxes matched in the first iteration and negative targets for other boxes and all negative classes. Note that we compute the parameter of the focal loss based on the frequency of positive targets per class in the current batch. Overall, the pathology detection loss is computed as .
Pathology Classification (Contrastive Loss)
We additionally apply a contrastive loss on the aggregated ROI token . Here, we consider the pathology targets , which define whether the pathology is present in the image. The ROI token for class is paired with its positive or negative pathology prompt or depending on . As negative pairs, we include the opposing prompt alongside all positive and negative prompts of all other pathologies. This leads to the following loss function:
where we set the temperature to .
0.D.2.2 Anatomy Tokens
Fig. 15 visualizes training with anatomy tokens. Given is the set of anatomical regions from CIG (29 regions from CIG and additionally 5 anatomical regions each consisting of 2 regions from CIG, e.g. lungs consisting of the right and left lungs). For each region , we encode its prompt using the prompt encoder and denote the resulting prompt tokens as . We pass these tokens to the prompt decoder, which predicts three ROI bounding boxes and the aggregated ROI token . Additionally, we encode positive and negative prompts for pathologies assigned to regions (we use 53 pathologies and findings from CIG) and denote the resulting prompt tokens as and , respectively.
Anatomy Detection (Bounding Box Loss)
For anatomical regions with a single target bounding box, we match all three predicted boxes for that region to the target bounding box. For regions with multiple target boxes, we apply the same matching strategy as used for pathology bounding boxes but using only as the matching cost. Finally, we apply the box loss to all matched bounding boxes. Box scores are not trained for anatomy tokens.
Anatomy Classification (Contrastive Loss)
We again apply a contrastive loss to the ROI tokens. Here we use the positive and negative pathology tokens and and align the ROI token of anatomy based on whether the pathologies are present in this region. The presence of pathology in anatomy is given as and is provided by CIG. This leads to the following loss function:
where we set the temperature to and randomly subsample a maximum of 10 negative classes (i.e. classes not present in any of the regions in the current batch).
Anatomy Sentence Generation
For sentence generation, we condition the sentence generator on the ROI tokens and train it to autoregressively predict the target sentences assigned to this region using the standard cross-entropy loss . Note that we concatenate sentences if multiple sentences are assigned to a region and ignore regions without sentences. Additionally, the MSE loss is applied between the output of the post decoder and the anatomy sentences encoded by the prompt encoder.
0.D.2.3 Sentence Tokens
Fig. 16 visualizes training with sentence tokens. Given are the sentences of sample from MIMIC-CXR. We encode each sentence using the prompt encoder and denote the resulting prompt tokens as . We pass each of these tokens to the prompt decoder, which predicts the aggregated ROI token for each of them.
Sentence Contrastive Loss
We again apply a contrastive loss between the ROI tokens and the prompt tokens . Unlike for pathology and anatomy tokens, where the contrastive loss was applied for each sample and token independently, we utilize sentence prompt tokens from the same and other samples in the batch as negatives. This leads to the following loss function:
where is the batch size and we set the temperature to .
Sentence Generation
For sentences generation, i.e. sentence reconstruction, we condition the sentence generator on the ROI tokens and train it to autoregressively reconstruct the report sentences used as prompts, using the standard cross-entropy loss . Additionally, the MSE loss is applied between the output of the post decoder and the encoded sentence prompts .
Global Contrastive Loss
To fine-tune and stabilize the image encoder, we additionally apply standard CLIP loss between the (global) average pooled projected patch features and the average pooled sentences tokens . We use as the temperature.
0.D.3 Multistage Training
We train ChEX in three different stages where we add new components with each stage while freezing some already trained components. Additionally, we prepare the language model for MIMIC-CXR reports in a text-only pre-training stage.
0.D.3.1 Stage 1
The goal of stage 1 is to train the prompt detector and fine-tune the image encoder. We initialize the image and text encoder as well as their projection heads from the CheXzero[70] weights of one of their checkpoints 111best_64_5e-05_original_22000_0.864.pt which we retrieved from their provided cloud drive222https://drive.google.com/drive/folders/1makFLiEMbSleYltaRxw81aBhEDMpVwno. The text encoder and its projection head are completely frozen throughout all training stages, and we freeze only the embedding layer and the first eight transformer layers of the image encoder for stage 1. The prompt detector is initialized randomly and will be trained in this stage. The sentence decoder is not yet used in this stage. We train using both datasets (MIMIC-CXR and VinDr-CXR) and all types of tokens – pathology tokens for samples from VinDr-CXR, anatomy tokens for the CIG targets for MIMIC-CXR samples, and sentence tokens for MIMIC-CXR samples. We use all bounding box and contrastive losses but do not yet train text generation. Overall, the loss function for stage 1 is defined as
Note that for anatomy tokens we only use the textual prompts in this stage, bounding box-based queries are not yet used. To improve generalization and limit the GPU memory consumption, we randomly select a subset from the possible tokens (i.e. queries for pathologies and anatomical regions) to be used in the current batch. We subsample 10 pathologies for bounding box training, 10 pathologies for the contrastive loss, and 20 anatomical regions.
We train with a batch size of 256 for 50 epochs on a single Nvidia A40, which takes around 36 hours. We train with AdamW[44] and use a cosine annealing schedule with a linear warmup of 1000 steps, initial learning rate and minimum learning rate Additionally, we use a weight decay of clip gradients with a norm larger than During training, we use the transformations from the Albumentations library [5]. We randomly crop and resize all images to a resolution of using a crop scale range of . We then apply random horizontal flips with probability, random affine transformations with rotation angle range , translation range , and scaling range . We also apply random contrast/brightness jittering with contrast/brightness range and random Gaussian blurring with probability and sigma range .
0.D.3.2 Stage 2
The goal of stage 2 is to train the post decoder of the sentence generator. We use the trained components from stage 1 but freeze all of them, i.e. we fully freeze the prompt encoder, image encoder, and prompt detector. We now add the (randomly initialized) post decoder of the sentence generator, however, without using the language model yet. We train only on the MIMIC-CXR dataset (including targets from CIG) and only use sentence and anatomy tokens. The VinDr-CXR dataset and pathology tokens are not used. We use the contrastive losses for anatomy and sentence tokens and additionally use MSE losses to reconstruct the sentence embeddings of target sentences based on ROI tokens from either anatomy or sentence prompts. We do not train bounding boxes or sentence generation. Overall, the loss function for stage 2 is defined as
For anatomical regions we randomly decide for each sample whether to use bounding-box-based or text-prompt-based queries. We train with a batch size of 256 and initial learning rate or 200 epochs on a single Nvidia A40, which takes around 24 hours. All other hyperparameters are kept the same as in stage 1.
0.D.3.3 Text-only Pre-Training
The goal of text-only pre-training is to prepare the decoder language model for MIMIC-CXR report sentences. We initialize the GPT-2 medium language model from a checkpoint333https://huggingface.co/healx/gpt-2-pubmed-medium pre-trained on PubMed abstracts. We only use the frozen prompt encoder together with the language model and a (randomly initialized) MLP projection for P-tuning, i.e. we do not use the post decoder. Using the sentences from MIMIC-CXR, we encode each sentence, apply the MLP projection, and reconstruct the original sentence using the language model. We train the language model and MLP projection using autoregressive training.
We use a batch size of reports (each with one or more sentences) but randomly drop sentences if there are more than 64 in a batch and use 8 gradient accumulation steps. We train with an initial learning rate of nd 10,000 warmup steps for 50 epochs on a single Nvidia A40, which takes around 3 days.
0.D.3.4 Stage 3
The goal of stage 3 is to train the whole sentence generator. We use the pre-trained post decoder from stage 2 and the decoder language model from text-only pre-training, both of which will be trained in this stage. Additionally, we use the prompt encoder, image encoder, and prompt detector from stage 1 but keep all of them frozen. We train only on the MIMIC-CXR dataset (including targets from CIG) and only use sentence and anatomy tokens. We autoregressive train the sentence generator to predict the target sentences of each anatomical region and to reconstruct the sentences for sentence tokens. Additionally, we use MSE losses between the output of the post decoder and the sentences encoded by the prompt encoder. Overall, the loss function for stage 3 is defined as
During training, we randomly drop the per-bounding-box ROI tokens for a query with a probability but always keep the aggregated ROI token. We use a batch size of reports (with gradient accumulation steps) but restrict the total number of generated sentences per batch to 32 for anatomical regions and 32 for sentence tokens, randomly subsampling if a batch contains more. We train with an initial learning rate of nd 10,000 warmup steps for 20 epochs on a single Nvidia A40, which takes around 4 days.
0.D.4 Inference
0.D.4.1 Sentence Grounding (SG) and Object Detection (OD)
For SG and OD, we use textual prompts as queries (the given input sentences in SG and predefined prompts, the names of pathology classes, for OD). The encoded prompts are given to the prompt detector, which predicts bounding boxes with box scores and ROI token vectors. For SG, we directly predict these bounding boxes with their associated box scores. For OD, we compute cosine similarities between the ROI tokens and positive as well as negative pathology prompts to compute box class probabilities. We then multiplicative combine these box class probabilities with the predicted box scores to get the final box scores. Finally, we apply task-specific post-processing as described in Sec. 0.E.2.
0.D.4.2 Region Classification (RC) and Explanation (RE)
For RC and RE, bounding boxes are given as inputs. We use these bounding boxes as inputs for the Gaussian ROI pooling layer of the prompt detector to compute ROI tokens for each input based on the patch features. Other parts of the prompt detector and the prompt encoder are not used for this task. We then further process the ROI tokens by the post decoder of the sentence decoder to further condition them on the patch features.
For RC, we now compute the cosine similarity between the post decoder outputs and encoded pathology prompts. Finally, we classify each region based on the cosine similarity. For the multiclass task MS-CXR, we compute the probability using softmax over all pathologies and predict the pathology class with maximum similarity. For the multilabel binary task CIG, we compute the softmax over the positive and negative prompts for each pathology class independently and predict all classes where the similarity with the positive pathology prompt is larger than the negative prompt.
For RE, we use the post decoder outputs to condition the language model and generate the descriptions for each region independently.
0.D.4.3 Full Report Generation (RG)
For RG, we use pre-defined sets of textual prompts. These prompts may include aspects like pathologies or anatomical regions as studied in Sec. 0.A.2. As pathology prompts, we use the set of 14 findings and pathologies from CheXpert[28]. As anatomy prompts we use a subset of 15 anatomical region names from the regions used in CIG, namely the lower, mid, and upper lung zones, hilar structures, apical zones, as well as the clavicle, each for left and right independently, and additional the mediastinum, the cardiac silhouette, and the abdomen. We encode the prompts with the prompt encoder and then apply the prompt detector to predict bounding boxes and ROI tokens for each prompt. The ROI tokens are then used by the sentence generator to predict textual descriptions for each prompt.
Next, we filter the descriptions to include in the final report. We compute the positive probability for each prompt (and the associated description) by classifying each ROI token similarly as in the region classification task using the CheXpert classes excluding no-finding. If any pathology is positive, we consider the prompt and associated description to be positive. We decided to keep all descriptions of pathology prompts but only positive anatomy descriptions (we refer to Sec. 0.A.2).
Finally, we filter the descriptions based on their associated box scores. We use a dynamic approach where we identify similar descriptions from different prompts by computing the cosine similarity between the descriptions encoded by the prompt encoder. For descriptions with high similarity, the lower-scoring ones are excluded. By setting the similarity threshold low enough, we allow the model to not only remove synonymous descriptions but also contradicting descriptions of similar aspects or several negative findings, where we only keep the most confident and relevant ones.
Appendix 0.E Evaluation Details
0.E.1 Benchmark Datasets
0.E.1.1 MIMIC-CXR
We use the MIMIC Chest X-ray (MIMIC-CXR) Database v2.0.0[30, 31, 14], which consists of chest X-ray images with corresponding free-text radiology reports from 227,835 radiographic studies of 65,079 patients performed at Beth Israel Deaconess Medical Center in Boston, US. The dataset authors de-identified the dataset satisfying the US Health Insurance Portability and Accountability Act of 1996 (HIPAA) Safe Harbor requirements and assuring Protected health information (PHI) has been removed. For all MIMIC-CXR-based datasets, i.e. MIMIC-CXR, CIG, and MS-CXR, we only use the frontal chest X-ray images from the already pre-processed MIMIC-CXR-JPG[32] version.
We download the free-text radiology reports from the original MIMIC-CXR dataset and extract the text from the Findings and Impression sections. Reports containing none of these sections, where both of these sections are empty, or which contain less than two words are removed. We split each of these sections into sentences using Stanza[57] and use all of these sentences (from both sections) as the report sentences of that sample. If a full report is required, as in the full report generation task, we concatenate all sentences to a single string. We follow the official MIMIC-CXR train/val/test split. This leads to a total of 1,879 images/report pairs for validation and 3,082 for testing.
0.E.1.2 Chest ImaGenome (CIG)
For anatomical regions, including bounding boxes with associated sentences and pathologies, we utilize the Chest ImaGenome dataset (CIG)[78, 79, 14]. It provides scene graphs automatically constructed from the free-text radiology reports from MIMIC-CXR. Each scene graph includes bounding boxes for 29 unique anatomical regions with associated attributes, where we consider 43 anatomical finding and 10 disease attributes as binary anatomy-level classification labels. Additionally, each region is associated with sentences from the report and we use these sentences as target sentences for anatomical regions. Note that not all sentences from the report may be associated with a region, sentences can be associated with several regions, and each region can be associated with zero, one, or multiple sentences. We only keep samples with scene graphs containing at least five valid region bounding boxes. We use the images from the MIMIC-CXR dataset (see above) and follow the official MIMIC-CXR train/val/test split444Note that this means that we do not follow the official split from CIG., leading to 1,959 images for validation and 3402 for testing. For multitask training of ChEX we combine the supervision from CIG with the sentences from MIMIC-CXR, leading to a total of 227,382 training images.
0.E.1.3 MS-CXR
The MS-CXR dataset [4, 3, 14], which we use only for evaluation, provides additional annotation for a subset of MIMIC-CXR. It contains 1,162 image–sentence pairs of bounding boxes and corresponding phrases for 8 pathologies annotated by board-certified radiologists. Each sentence is associated with a pathology and one or more bounding boxes. Images in MS-CXR contain at least one sentence (and pathology) but may also contain several sentences with possibly different pathologies. We use the images from the MIMIC-CXR dataset. However, only very few images of MS-CXR are part of the MIMIC-CXR test split. For MS-CXR tasks, we thus use all MS-CXR images from the validation and test split of MIMIC-CXR for testing but exclude all images from the training split to avoid any data leakage. This leads to a total of 169 test images.
0.E.1.4 VinDr-CXR
We use the VinDr-CXr[54, 53, 14] dataset, containing 18,000 postero-anterior (PA) view, i.e. frontal, chest X-ray images retrospectively collected from the Hospital 108 (H108) and the Hanoi Medical University Hospital (HMUH), two of the largest hospitals in Vietnam. The images were manually annotated by radiologists with 22 localized labels (with bounding boxes) and 6 global (image-level classification) labels of diseases (pathologies). Each image may be associated with zero, one, or more pathologies, and for each (localized) pathology there may be one or more provided bounding boxes. In order to protect the patient’s privacy and meet the U.S. HIPAA, the European GDPR, and the local privacy laws, all personally identifiable information associated with the images has been removed or replaced with random values[53]. The images are provided in the DICOM format, and we converted them into PNG files using the pydicom555https://pydicom.github.io/ library. The dataset authors provide a train/test split consisting of 15,000 and 3,000 samples, respectively. The samples in the train split contain bounding box annotations from individual radiologists that have not been merged by the dataset authors. We fuse bounding box annotations from different radiologists corresponding to the same pathology using weighted box fusion[66] with IoU threshold . Samples from the provided test split come with merged annotations, such that no further processing is required. We use all samples from the training split for training and randomly split the 3,000 test split samples into 1,500 samples for validation and 1,500 samples for testing.
0.E.1.5 NIH ChestXray8 (NIH8)
We use the ChestXray-8 (NIH8) dataset[75], which consists of 108,948 chest X-ray images retrospectively collected from the National Institutes of Health Clinical Center in the US. The authors do not provide information about data anonymization. The images are provided as pre-processed PNG-images (extracted from DICOM files) such that no further pre-processing is applied.
The dataset contains a subset of images with bounding boxes for eight different disease types hand-labeled by a board-certified radiologist. We randomly split these images into 432 samples for validation and 448 samples for testing.
0.E.2 Benchmark Task Evaluation
0.E.2.1 Task-specific Inference Hyperparameters
While no task-specific fine-tuning is applied, we use task-specific inference and post-processing hyperparameters for ChEX and all baselines to consider differences in annotation practices of datasets. We tune those hyperparameters on samples from the validation splits of the corresponding datasets.
For the sentence grounding (SG) task, we scale all predicted boxes by a constant factor which we optimize on the validation set. Additionally, we apply NMS with an IoU threshold of . For the pathology detection (OD) tasks, we scale boxes independently for each pathology class. We apply NMS with an IoU threshold of for the VinDr-CXR and MS-CXR datasets. For the NIH8 dataset, we merge all predicted boxes per pathology (using the super bounding box covering all predicted boxes) and report the maximum of the predicted score as its box score.
0.E.2.2 Metrics and Evaluation
For SG, we use the sentences from MS-CXR as input and the corresponding bounding boxes as targets. We compute mIoU using the thresholds . For each sentence and threshold, we select the predicted bounding boxes based on the box scores and merge them into a single mask (corresponding to one or more boxes). Similarly, we merge the target boxes of that sentence before computing the IoU between the predicted and target mask. Finally, the mIoU is computed by first averaging the IoU per threshold over all sentences before averaging over all thresholds. For mAP computation, we treat the task as object detection, where each sentence corresponds to its own class, and then use the IoU thresholds .
For OD, we use the pathology classes provided in the three datasets (VinDR-CXR, MS-CXR, NIH8), where for VinDr-CXR, we only use the top 15 most frequent classes. We report the standard mAP metric with IoU thresholds .
For RC, we provide bounding boxes of regions (pathologies for MS-CXR, anatomical regions for CIG) as input and expect the model to classify these regions. On MS-CXR this is a multiclass classification problem over 8 classes. Thus, we use the macro-averaged AUROC metric (one-vs-rest). On CIG, this is a multilabel binary classification problem with 53 highly imbalanced classes. Thus, we use the weighted AUROC (wAUROC) metric, which considers the frequency of classes.
For RE tasks, we provide bounding boxes of regions (pathologies for MS-CXR, anatomical regions for CIG) as input and expect the model to predict textual descriptions for them. On the CIG dataset, regions without associated sentences are ignored, and for regions with multiple sentences, we concatenate their sentences. For metric computation, we consider each region as its individual sample.
For RG, we only provide the image and expect the model to predict the full report, where each image/report pair is considered a single sample. Baseline results are taken from their publications. Note that some baselines have been evaluated on different test splits and that pre- or post-processing can differ as well, leading to limitations in the interpretation of exact results as also acknowledged by [26, 69].
Final results are reported based on samples from the test split. We use bootstrapping with for the SG, OD, and RC tasks while using for the RE and RG tasks (due to expansive metric computation).
0.E.3 Evaluation of Interactive Capabilities
We evaluated the interactive capabilities as studied in Secs. 5.2 and 5.3 on a small subset of the test sets of CIG and MS-CXR. We use only frontal images that have both CIG and MS-CXR targets and exclude samples where the positive classes present in the MS-CXR targets are not present in the CIG targets, i.e., where CIG and MS-CXR targets are not consistent. This results in a total of 98 images.
[…]
For Fig. 6b, we apply the CheXbert classifier on predicted sentences and identify the number of positive predicted pathology classes. We refer to the pathology used in the prompt as queried and to all other pathologies as non-queried.
0.E.4 Baseline Details
0.E.4.1 Supervised Visual Grounding Baselines
We train the supervised visual grounding baselines on the 878 MS-CXR samples corresponding to the MIMIC-CXR training split. For TransVG[9] we use their official implementation666https://github.com/djiajunustc/TransVG. We additionally study modified versions of TransVG, where we replace its default image backbone (ResNet50[19]) with backbones pre-trained on chest X-ray images (MIMIC-CXR). Specifically, we use the image encoders from BioVIL[4] and CheXzero[70]. BioVIL provides a pre-trained ResNet50 backbone, such that the default backbone can be initialized with BioVIL weights without requiring further modifications. CheXzero only provides a ViT backbone, such that the whole backbone is completely replaced by the CheXzero backbone. Here, we use the patch tokens of the final transformer layer to replace the patch features of the ResNet.
Chen et al. [7] proposed a supervised visual grounding model for the MS-CXR dataset. However, their model shows only marginal improvements above TransVG, and no public source code is provided, such that we did not include it in our studies. Also, they evaluated their model only on samples with a single bounding box per sentence, such that their results are not comparable with our benchmark.
0.E.4.2 Supervised Object Detection Baselines
We train all supervised object detection baselines on the VinDr-CXR dataset. For evaluation on other OD tasks, we map the VinDr-CXR pathology classes to the corresponding classes of the downstream task. With this strategy, we managed to map all classes of MS-CXR and NIH8, except for pneumonia.
For Faster R-CNN[63] we use the torchvision777https://pytorch.org/vision implementation. For DETR[6], Conditional DETR[47], and Deformable DETR[90] we use the implementations from the Huggingface transformers library888https://huggingface.co/docs/transformers. For o-Deformable-DETR[91] we use the implementation provided by the MMDetection library999https://mmdetection.readthedocs.io. We again study modified versions of these models, where we replace their default backbones with BioVIL[4] and CheXzero[70], following the same approach as for TransVG. Note that we do not study the CheXzero backbone with advanced DETR-style models (Conditional DETR, Deformable DETR, and Deformable DETR) as the standard DETR model with the CheXzero backbone performs very poorly.
We additionally evaluate the Faster R-CNN models on the region classification (RC) tasks. Here, we modify their ROI pooling layer and replace the predicted box proposals with the bounding boxes provided as input. With this modification, their ROI classification heads can be used to classify the given regions. Again, VinDr-CXR classes are mapped to MS-CXR and CIG classes, where we manage to map 7 of 8 MS-CXR classes and 17 of 53 (including many of the most frequent) CIG classes. Note that DETR-based models cannot be utilized for region classification as they do not have an ROI pooling layer.
0.E.4.3 Weakly-Supervised Object Detection Baselines
We also compare against weakly-supervised object detection (WSupOD) models. Such models use only limited supervision without bounding boxes for pathologies during training but can predict pathology bounding boxes during inference.
We study CheXnet[60], a model commonly used on chest X-rays. We use our own implementation with the DenseNet121[22] backbone and train it on the 91,883 NIH8 training samples using only image-level classification targets. Additionally, we again replace the default backbone with BioVIL[4] and CheXzero[70]. For evaluation on other datasets, we map the NIH8 pathologies to the corresponding classes of the downstream task, where we managed to map all MS-CXR classes, 7 of the 15 VinDr-CXR classes used for evaluation, and 10 of the 53 CIG classes. We also evaluate the CheXnet models on the region classification (RC) tasks by replacing the average-pooling layer before the classifier by ROI pooling with the bounding box provided for the region.
We also study the two models MIL-ADPD[51] and Loc-ADPD[51]. Both use anatomical region bounding boxes and pathology classification targets, where MIL-ADPD uses image-level classification targets while Loc-ADPD uses region-level targets. We use their official implementation101010https://github.com/philip-mueller/adpd and trained it on the 237,917 samples from the CIG training dataset (i.e. the CIG samples from the MIMIC-CXR training split). We again replace their default backbone with BioVIL[4] but do not study the use of the CheXzero backbone as it performs very poorly on WSupOD with CheXnet. We again mapped pathology classes from CIG to the target task, where we were able to successfully map all classes.
0.E.4.4 Contrastive Baselines
We study two contrastive models pre-trained on the MIMIC-CXR dataset, namely BioVIL[4] and CheXzero[70]. We use the official implementations and checkpoints for both BioVIL111111https://github.com/microsoft/hi-ml/tree/main/hi-ml-multimodal and CheXzero121212https://github.com/rajpurkarlab/CheXzero, where for CheXzero we use only one of their checkpoints131313best_64_5e-05_original_22000_0.864.pt which we retrieved from their provided cloud drive141414https://drive.google.com/drive/folders/1makFLiEMbSleYltaRxw81aBhEDMpVwno.
For sentence grounding, we follow the approach of Boecking et al. [4]. We first encode the input sentences using their text encoder and the patch features using their image encoder. Next, we project both into the shared image-text space and compute the cosine similarity between the sentence and each patch. The resulting similarity map is interpolated to the image resolution and then thresholded to generate masks, which are then used to compute the mIoU with the target bounding boxes. Additionally, we apply the thresholding procedure from CheXnet[60] on the similarity maps to predict bounding boxes. We compute box scores based on the (average) similarity map values overlapping with the bounding box. The box prediction enables a better comparison with box-based models and allows the computation of both mIoU and mAP scores.
For object detection, we follow a similar approach to predict bounding boxes but use predefined textual prompts for pathologies instead of sentences.
For region classification, we first compute embeddings for textual prompts of each of the classes. We then compute image features for the given region bounding boxes and compute the cosine similarity between the projected image and text features. For the multiclass task on MS-CXR, we apply softmax over the similarities of all classes, for the multilabel task on CIG, we use positive and negative prompts and apply softmax over the positive and negative similarities for each class independently. We experiment with two approaches to compute the region image features, ROI pooling and cropping. In the ROI pooling approach, we encode the whole image and then simply pool the patch features overlapping with the given region bounding boxes. For cropping, we feed each region as its individual sample, cropping the image to the given region.
For region explanation, we utilize the image-text retrieval capabilities of the contrastive models. We encode all sentences from the training set independently and store them in a vector database. During inference, we encode each region using the ROI pooling approach from region classification and find the sentence with the highest cosine similarity.
0.E.4.5 Generative Baselines
We evaluate the RGRG[69] model for region explanation using their official implementation and checkpoint151515https://github.com/ttanida/rgrg. For full report generation on MIMIC-CXR, we report the values from the original publications.